What is the nth term for this pattern?....... 2,8,18,32,50
Found 2 solutions by scott8148, issacodegard:
Answer by scott8148(6628)

dividing by 2 ___ 1, 4, 9, 16, 25 ___ look familiar?

Answer by issacodegard(60)

Hi, the first term is 2, and I'll denote the nth term by f(n). The sequence of terms for f(n)/2 is 1,4,9,16,25,... which are the positive integers squares. So we have f(n)/2=n^2. Therefore, f(n)=2n^2.